If a consumer has a utility function u(x1, x2) = x1x24 , what fraction of her income will be spent on good 2?
Answer #1

Answer

u(x1, x2) = x1x24

This function is a cobb douglas utility function and according to properties of a cobb douglas function, If a consumption function is of the form U(x1, x2) = Ax1ax2b, then consumer will spend a/(a + b) portion of Income on Good x1 and b/(a + b) fraction of income on good x2.

Hence for this question, a = 1 and b = 4 fraction of Income spend on x2 is b/(a + b) = 4/(1 + 4) = 4/5

Hence, fraction of her income that will be spent on good 2 is 4/5